• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Camsonne

    Log entry time 00:40:13 on November 8,2004

    Entry number 136116

    This entry is a followup to: 136063

    keyword=Start of run : forgot update settings

    I noticed that the data for the spectrometer have not been updated :
    Left spectrometer is at 2.95 Gev/c and 19.32 degrees